# Stopping Rules for Stochastic Gradient Descent via Anytime-Valid Confidence Sequences

### v3

@@ 2 Anytime-Valid Confidence Sequences
In this subsection, we unpack the proof of Theorem 2.1 through a sequence of conceptual steps. Rather than presenting all technical details at once, we isolate the key ideas that transform the projected SGD recursion into an anytime-valid confidence sequence. We adopt this layered presentation, along with detailed explanations, deliberately for its didactic value. We first construct a process whose one-step evolution compares the actual suboptimality to what is explained by the geometry of projected SGD and the stochastic gradients, and show that this process enjoys a trajectory-dependent sub-Gaussian control. We then present a time-uniform concentration lemma for martingale sums, after which a simple telescoping argument and normalization by $S_{t}$ yield the desired bound on $\bar{F}_{t}$.
@@
with $X_{0}\coloneqq 0$. The next lemma shows that $X_{t}$ has two crucial properties: it has nonpositive conditional drift, and its conditional moment generating function is dominated by that of a sub-Gaussian variable with variance proxy $\sigma_{t}^{2}=\sigma^{2}\eta_{t}^{2}Z_{t}$, as defined in (8). This is where convexity of $f$, the unbiasedness assumption, and the conditional sub-Gaussian noise assumption enter.
@@
which proves (14).
@@
Applying Assumption 1.1(iii) with
@@
To state the result, define
$c_{1}\;:=\;\max\!\left\{1,\;\frac{\log\Big(\log\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}+1\Big)}{\log\Big(\log(\mathrm{e}+1)\Big)}\right\},$ (20)
with $\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}$ defined in (11), and set $\lambda_{0}:=\sqrt{c_{1}/\sqrt{\mathrm{e}}}$. Moreover, for $k=0,1,2,\ldots$, define
$\lambda_{k}:=\lambda_{0}\mathrm{e}^{-k/2},\qquad w_{k}:=\frac{6}{\pi^{2}(k+1)^{2}}.$
@@
$\forall t\geq 1:\ \bar{X}_{t}\leq C\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\Big(\log\frac{2}{\alpha}+\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})\Big)}.$
with constant $C=6\sqrt{c_{1}}$ and $c_{1}$ defined in (20).
@@
Using $w_{k}={6}/(\pi^{2}(k+1)^{2})$ and $\lambda_{k}=\lambda_{0}\mathrm{e}^{-k/2}$, we have
@@
and
$\frac{1}{\lambda_{k}}=\frac{\mathrm{e}^{k/2}}{\lambda_{0}}.$
Hence, for $v>0$,
$B_{k}(v):=\frac{\log(1/(\alpha w_{k}))}{\lambda_{k}}+2\lambda_{k}v\leq\frac{\mathrm{e}^{k/2}}{\lambda_{0}}\big(L_{\alpha}+2\log(k+1)\big)+2\lambda_{0}\mathrm{e}^{-k/2}v.$
@@
which immediately implies that $\mathrm{e}^{k(v)/2}\leq z_{v}\leq\mathrm{e}^{(k(v)+1)/2}$. Since $\Sigma_{t,\mathrm{eff}}^{2}\geq 2(\log\tfrac{2}{\alpha}+1)$ by definition, Lemma A.3 guarantees that $k(v)\geq 0$. For the lemma, we note that $L_{\alpha}\geq 1$ holds since $\alpha\leq\tfrac{2}{\mathrm{e}}$. Therefore, for this choice, we have that
$\frac{1}{\sqrt{\mathrm{e}}}\,z_{v}\leq\mathrm{e}^{k(v)/2}\leq z_{v},\qquad\log(k(v)+1)\leq c_{1}L_{v},$ (22)
with the constant $c_{1}\geq 1$ defined in (20). Note that the inequality $\log(k(v)+1)\leq c_{1}L_{v}$ follows from Lemma A.2, using the fact that $L_{\alpha}+L_{v}>L_{\alpha}$. For this $k=k(v)$, let
$D_{v}:=L_{\alpha}+2\log(k(v)+1).$
Using (22), we have $2\log(k(v)+1)\leq 2c_{1}L_{v}$, so
$D_{v}\leq L_{\alpha}+2c_{1}L_{v}\leq c_{2}\big(L_{\alpha}+L_{v}\big)$
for the constant $c_{2}\coloneqq\max\{1,2c_{1}\}=2c_{1}$.
@@
$\frac{\mathrm{e}^{k(v)/2}}{\lambda_{0}}D_{v}\leq\frac{1}{\lambda_{0}}\sqrt{\frac{v}{L_{\alpha}+L_{v}}}\,c_{2}\big(L_{\alpha}+L_{v}\big)=\frac{c_{2}}{\lambda_{0}}\sqrt{v\big(L_{\alpha}+L_{v}\big)},$
@@
$2\lambda_{0}\mathrm{e}^{-k(v)/2}v\leq 2\lambda_{0}\sqrt{\mathrm{e}}\sqrt{\frac{L_{\alpha}+L_{v}}{v}}\,v=2\sqrt{\mathrm{e}}\lambda_{0}\sqrt{v\big(L_{\alpha}+L_{v}\big)}.$
Summing these two bounds and recalling that $\lambda_{0}=\sqrt{{c_{2}}/({2\sqrt{\mathrm{e}}})}$, we obtain
$B_{k(v)}(v)\leq 2\sqrt{2\sqrt{\mathrm{e}}c_{2}}\sqrt{v\big(L_{\alpha}+L_{v}\big)}\leq c_{3}\sqrt{v\big(L_{\alpha}+L_{v}\big)},$
for the constant $c_{3}\coloneqq 4\sqrt{c_{2}}\leq 6\sqrt{c_{1}}=:C$.
@@
$\inf_{k\geq 0}B_{k}(v)\leq C\sqrt{v\big(L_{\alpha}+\log\log(\mathrm{e}+v)\big)}=C\sqrt{v\big(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+v)\big)}.$
@@
$\bar{X}_{t}\leq C\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\Big(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})\Big)}\qquad\forall t\geq 1,$
@@
$\sum_{s=1}^{t}\eta_{s}\big(f(x_{s})-f(x^{\star})\big)\;\leq\;\frac{1}{2}\Bigg(C\,\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\Big(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})\Big)}+Z_{1}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}\Bigg),$ (23)
where $C>0$ is the constant from Lemma 2.6.
@@
$\forall t\geq 1:\qquad\sum_{s=1}^{t}X_{s}\leq C\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\Big(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})\Big)}.$ (24)
@@
$2\sum_{s=1}^{t}\eta_{s}\big(f(x_{s})-f(x^{\star})\big)\leq C\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\Big(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})\Big)}+Z_{1}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}.$
@@
This lemma is the bridge from the martingale world back to optimization: it translates the time-uniform control of the process $\bar{X}_{t}$ into a time-uniform upper bound on the cumulative weighted suboptimality $\sum_{s=1}^{t}\eta_{s}(f(x_{s})-f^{\star})$. Finally, dividing both sides by $S_{t}$ yields a time-uniform bound for the weighted average suboptimality $\bar{F}_{t}$, completing the proof of Theorem 2.1.
@@ 2.2 Clarifying remarks and intuition
The proof of Theorem 2.1 relies on several ideas from time-uniform sequential analysis that may be unfamiliar to readers with a background in stochastic optimization. In this subsection we collect three clarifying remarks intended to illuminate the construction. First, we explain the role of the mixture supermartingale $\bar{E}_{t}$ and why a single-parameter exponential martingale cannot yield a time-uniform bound without sacrificing optional stopping. Second, we provide a heuristic “continuous optimization” argument that motivates the geometric grid over $\lambda_{k}$ and the emergence of an iterated-logarithm term in the boundary. Third, we discuss why our bound does not contain a $\log\log t$ factor typical in classical anytime concentration results, and how the square-summable stepsize condition fundamentally alters the variance accumulation in the stochastic approximation setting.
@@
The mixture construction in Lemma 2.6 reconciles these goals by fixing a countable family of parameters a priori and combining the corresponding exponential supermartingales. This “method of mixtures” (also called pseudo-maximization) goes back to Robbins and Siegmund’s classical work on sequential boundary crossing and is developed systematically in the theory of self-normalized processes (see [de2009self]). Modern confidence-sequence constructions build on this principle in various forms, including explicit mixture and stitching frameworks such as the sub-$\psi$ approach in [howard2021time], as well as betting- and e-process-based methods that aggregate families of nonnegative supermartingales; see [waudby2024estimating]. $\clubsuit$
@@
The mixture construction in Lemma 2.6 hinges on choosing a geometric grid of tuning parameters $\lambda_{k}=\lambda_{0}\mathrm{e}^{-k/2}$. Here we provide intuition for this choice by examining how the envelope $B_{k}(\Sigma_{t,\mathrm{eff}}^{2})$ is minimized, and how the geometry of the grid matches the natural scale of the problem. Let $v\coloneqq\Sigma_{t,\mathrm{eff}}^{2}$ and recall the upper envelope
$B_{k}(v):=\frac{\log(1/(\alpha w_{k}))}{\lambda_{k}}+2\lambda_{k}v,\qquad\lambda_{k}=\lambda_{0}\mathrm{e}^{-k/2},\quad w_{k}=\frac{6}{\pi^{2}(k+1)^{2}}.$
@@
$z^{\star}=\sqrt{\frac{2\lambda_{0}^{2}v}{D}}\ \implies\ \lambda^{\star}=\frac{\lambda_{0}}{z^{\star}}=\sqrt{\frac{D}{2v}}.$
@@
The slowly varying term $\log(k+1)$ introduces an iterated-logarithm correction, where the $\log\log(\mathrm{e}+v)$ factor arises precisely from the slow growth of $\log(k+1)$ on this geometric grid. In particular, for a fixed time $t$, the optimal value of $\lambda$ is of order $v^{-1/2}$, where $v=\Sigma_{t,\mathrm{eff}}^{2}$ controls the variance scale of the martingale sum. Since the scale $\Sigma_{t,\mathrm{eff}}^{2}$ is not known in advance and may vary with the trajectory up to the a priori bound $\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}$, a geometric grid of the form $\lambda_{k}=\lambda_{0}\mathrm{e}^{-k/2}$, with $\lambda_{0}$ chosen as a function of $\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}$, is a natural discretization: consecutive grid points differ by a constant factor, ensuring that for every $t$ there exists a $\lambda_{k}$ within a constant factor of the ideal choice $\lambda^{\star}\approx\Sigma_{t,\mathrm{eff}}^{-1}$.
@@
In short, the grid is chosen so that: (i) it spans the relevant scales of $\Sigma_{t,\mathrm{eff}}^{2}$ via exponentially spaced search points, and (ii) at least one $\lambda_{k}$ is guaranteed to be within a constant factor of the ideal minimizer, thereby achieving near-optimal behavior across all times $t$. This “scale-matching” intuition for geometric discretizations is common in mixture and stitching arguments for martingale concentration; see, for example, the heuristic discussion surrounding Theorem 1 in [howard2021time] and the broader treatment of pseudo-maximization via mixtures in self-normalized processes [de2009self]. $\clubsuit$
Remark 2.10 (Why no $\log\log(\mathrm{e}+t)$ term appears).
A recurring feature in time-uniform martingale concentration is the appearance of iterated logarithm corrections, such as $\log\log(\mathrm{e}+t)$; see, for example, classical boundary-crossing analyses of Robbins and Siegmund [robbins1970boundary] and modern confidence-sequence treatments such as [howard2021time]. At first glance, one might therefore expect a confidence sequence for stochastic optimization to include a term of the form $\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\log\log(\mathrm{e}+t)}$ rather than the $\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})}$ term appearing in Theorem 2.1. The difference comes from the square-summable stepsize condition in Assumption 1.2 and the resulting definition of $\Sigma_{t,\mathrm{eff}}^{2}$. In most sequential testing settings, $\Sigma_{t,\mathrm{eff}}^{2}$ grows linearly with $t$, so controlling $k$ in Step (iv) of Lemma 2.6 ultimately leads to iterated-logarithm terms in $\log(\mathrm{e}+t)$. Here, however, $\{\eta_{t}\}_{t\geq 1}$ is square summable, so $\Sigma_{t,\mathrm{eff}}^{2}$ is uniformly bounded by $\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}<\infty$. Consequently, the relevant “scale” for the confidence sequence is the variance process itself rather than the time index. This variance-based iterated logarithm therefore remains bounded and is substantially smaller than the classical time-based correction that arises when the variance grows linearly with $t$.
This phenomenon is well known in the martingale boundary literature: confidence sequences expressed in terms of data-dependent variance proxies (such as predictable quadratic variation, conditional variance processes, or self-normalized sums) adapt to the realized variability of the process and can be substantially tighter when the variance grows slowly or remains bounded, compared to bounds written purely in terms of time. See, for example, [de2009self] and the discussion in [howard2021time, Section 3] on expressing time-uniform bounds in terms of intrinsic variance processes rather than the time index. $\clubsuit$
3 Anytime-Valid Stopping Rules
A central motivation for our confidence-sequence approach is to turn projected SGD into an anytime-valid optimization procedure: instead of committing to a fixed horizon in advance, the user would like to stop the algorithm based on the information actually observed along the trajectory. In practice, most stopping rules for SGD are heuristic (i.e., they monitor validation loss, gradient norms, or empirical convergence patterns) and their guarantees, when available, typically hold only for a pre-specified time or for a specific deterministic stopping rule. In contrast, a confidence sequence is designed precisely to support optional stopping: it provides a high-probability guarantee that remains valid under arbitrary data-dependent monitoring and stopping strategies.
Corollary 2.4 provides an observable anytime-valid upper confidence sequence $\{U_{t}^{\mathrm{obs}}(\alpha)\}_{t\geq 1}$ for the weighted average suboptimality $\bar{F}_{t}$ in (4). Combining this with the convexity bound $f(\bar{x}_{t})-f^{\star}\leq\bar{F}_{t}$, we obtain, for any fixed $\alpha\in(0,1)$,
$\mathbb{P}\!\left(\forall t\geq 1:\ f(\bar{x}_{t})-f^{\star}\leq U_{t}^{\mathrm{obs}}(\alpha)\right)\;\geq\;1-\alpha,$ (31)
Given a target accuracy $\varepsilon>0$ and confidence level $\alpha\in(0,\tfrac{2}{\mathrm{e}})$, it is therefore natural to stop as soon as the certificate $U_{t}^{\mathrm{obs}}(\alpha)$ falls below $\varepsilon$. As already foreshadowed in the introduction, we formalize this through the stopping time
$\tau_{\varepsilon}:=\inf\big\{t\geq 1:U_{t}^{\mathrm{obs}}(\alpha)\leq\varepsilon\big\},$ (32)
with the convention $\inf\emptyset=\infty$. Since $U_{t}^{\mathrm{obs}}(\alpha)$ is $\mathcal{F}_{t}$-measurable for each $t$, $\tau_{\varepsilon}$ is a proper stopping time with respect to the filtration $\{\mathcal{F}_{t}\}_{t\geq 1}$ generated by the iterates and gradients.
In the remainder of this section we show that, under the stochastic approximation conditions on the stepsize in Assumption 1.2, the stopping time $\tau_{\varepsilon}$ enjoys two key properties: (i) it is almost surely finite for any fixed $\varepsilon>0$, and (ii) the returned iterate $\bar{x}_{\tau_{\varepsilon}}$ is $\varepsilon$-optimal with confidence $1-\alpha$. Together, these facts certify projected SGD as an anytime-valid $\varepsilon$-optimization algorithm driven entirely by observable quantities along the run.
Theorem 3.1 (Adaptive anytime-valid stopping rule).
Suppose Assumptions 1.1 and 1.2 hold, and let $\{U_{t}^{\mathrm{obs}}(\alpha)\}_{t\geq 1}$ be the anytime-valid upper confidence sequence from Corollary 2.4. Fix $\varepsilon>0$ and $\alpha\in(0,\tfrac{2}{\mathrm{e}})$, and define the stopping time $\tau_{\varepsilon}$ by (32). Then:
  1. (i)
    $\tau_{\varepsilon}<\infty$ almost surely;
  2. (ii)
    the stopped iterate $\bar{x}_{\tau_{\varepsilon}}$ satisfies:
    $\mathbb{P}\Big(f(\bar{x}_{\tau_{\varepsilon}})-f^{\star}\leq\varepsilon\Big)\;\geq\;1-\alpha.$ (33)
In particular, projected SGD equipped with $\tau_{\varepsilon}$ yields a data-dependent stopping strategy that guarantees $\varepsilon$-optimality with probability $1-\alpha$, without requiring a predetermined time horizon.

### v4

@@ 3.1 Stopping-time complexity
We prove Theorem 2.1 by decomposing the projected SGD recursion into a sequence of conceptual steps. We first construct a one-step process that compares the instantaneous suboptimality to the geometry of projected SGD and the stochastic gradient noise, and show that this process satisfies a trajectory-dependent sub-Gaussian bound. We then derive a time-uniform concentration bound for the resulting martingale sum, after which a simple telescoping argument and normalization by $S_{t}$ yield an anytime-valid upper confidence sequence for the weighted suboptimality $\bar{F}_{t}$.
@@
with $X_{0}\coloneqq 0$. The next lemma shows that $X_{t}$ has two crucial properties: it has nonpositive conditional drift, and its conditional moment generating function is dominated by that of a sub-Gaussian variable with variance proxy on the order of $\sigma_{t}^{2}=\sigma^{2}\eta_{t}^{2}Z_{t}$, as defined in (5). This is where convexity of $f$, the unbiasedness assumption, and the conditional sub-Gaussian noise assumption enter.
@@
which proves (16). Note that $x_{t}$ is $\mathcal{F}_{t-1}$-measurable for each $t\geq 2$, since it is a deterministic function of $(x_{1},g_{1},\ldots,x_{t-1},g_{t-1})$ via (1).
@@
Applying Assumption 1.2(ii) with
@@
To state the result, define, for $k=0,1,2,\ldots$,
$\lambda_{k}:=\mathrm{e}^{-k/2},\qquad w_{k}:=\frac{6}{\pi^{2}(k+1)^{2}}.$
@@
$\forall t\geq 1:\ \bar{X}_{t}\leq 7\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\Big(\log\frac{2}{\alpha}+\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})\Big)}.$
@@
Using $w_{k}={6}/(\pi^{2}(k+1)^{2})$ and $\lambda_{k}=\mathrm{e}^{-k/2}$, we have
@@
and $1/{\lambda_{k}}=\mathrm{e}^{k/2}$. Hence, for $v>0$,
$B_{k}(v):=\frac{\log(1/(\alpha w_{k}))}{\lambda_{k}}+2\lambda_{k}v\leq\mathrm{e}^{k/2}\big(L_{\alpha}+2\log(k+1)\big)+2\mathrm{e}^{-k/2}v.$
@@
which immediately implies that $\mathrm{e}^{k(v)/2}\leq z_{v}\leq\mathrm{e}^{(k(v)+1)/2}$. Since $v=\Sigma_{t,\mathrm{eff}}^{2}\geq 2(\log\tfrac{2}{\alpha}+1)$ by definition, Lemma B.4 guarantees that $k(v)\geq 0$. For the lemma, we note that $L_{\alpha}\geq 1$ holds since $\alpha\leq\tfrac{2}{\mathrm{e}}$. Therefore, for this choice, we have that
$\frac{1}{\sqrt{\mathrm{e}}}\,z_{v}\leq\mathrm{e}^{k(v)/2}\leq z_{v},\qquad\log(k(v)+1)\leq L_{v}+1,$ (23)
where the second inequality in (23) can be justified as follows:
$\log\!\big(k(v)+1\big)\;\leq\;\log\!\left(\log\frac{v}{L_{\alpha}+L_{v}}+1\right)\;\leq\;\log(\log v+1)\;\leq\;\log(\mathrm{e}\log v)\;\leq\;1+\log\log(\mathrm{e}+v).$
Here, the first inequality follows from the definition of $k(v)$. The second inequality uses the fact that $L_{\alpha}+L_{v}>1$. The third inequality follows from $\log v+1\leq\mathrm{e}\log v$ for all $v\geq\mathrm{e}$. Finally, the last inequality holds because $\log(\mathrm{e}+v)\geq\log v$. For this choice of $k=k(v)$, define
$D_{v}:=L_{\alpha}+2\log\big(k(v)+1\big).$
Using (23), we have $2\log(k(v)+1)\leq 2L_{v}+2$, and therefore
$D_{v}\leq L_{\alpha}+2+2L_{v}\leq 3\big(L_{\alpha}+L_{v}\big),$
where the last inequality follows from the fact that $L_{\alpha}\geq 1$.
@@
${\mathrm{e}^{k(v)/2}}D_{v}\leq 3\sqrt{v\big(L_{\alpha}+L_{v}\big)},$
@@
$2\mathrm{e}^{-k(v)/2}v\leq 2\sqrt{\mathrm{e}}\sqrt{\frac{L_{\alpha}+L_{v}}{v}}\,v=2\sqrt{\mathrm{e}}\sqrt{v\big(L_{\alpha}+L_{v}\big)}.$
Summing these two bounds, we obtain
$B_{k(v)}(v)\leq\big(3+2\sqrt{\mathrm{e}}\big)\sqrt{v\big(L_{\alpha}+L_{v}\big)}\leq 7\sqrt{v\big(L_{\alpha}+L_{v}\big)}.$
@@
$\inf_{k\geq 0}B_{k}(v)\leq 7\sqrt{v\big(L_{\alpha}+\log\log(\mathrm{e}+v)\big)}=7\sqrt{v\big(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+v)\big)}.$
@@
$\bar{X}_{t}\leq 7\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\Big(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})\Big)}\qquad\forall t\geq 1,$
@@
$\sum_{s=1}^{t}\eta_{s}\big(f(x_{s})-f(x^{\star})\big)\;\leq\;\frac{1}{2}\Bigg(7\,\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\Big(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})\Big)}+Z_{1}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}\Bigg).$ (24)
@@
$\forall t\geq 1:\qquad\sum_{s=1}^{t}X_{s}\leq 7\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\Big(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})\Big)}.$ (25)
@@
$2\sum_{s=1}^{t}\eta_{s}\big(f(x_{s})-f(x^{\star})\big)\leq 7\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\Big(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})\Big)}+Z_{1}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}.$
@@
This lemma is the bridge from the martingale world back to optimization: it translates the time-uniform control of the process $\bar{X}_{t}$ into a time-uniform upper bound on the cumulative weighted suboptimality $\sum_{s=1}^{t}\eta_{s}(f(x_{s})-f(x^{\star}))$. Finally, dividing both sides by $S_{t}$ yields a time-uniform bound for the weighted average suboptimality $\bar{F}_{t}$, completing the proof of Theorem 2.1.
@@ 4.2 Clarifying remarks and intuition
The proof of Theorem 2.1 relies on several ideas from time-uniform sequential analysis that may be unfamiliar to readers with a background in stochastic optimization. In this subsection we collect two clarifying remarks intended to illuminate the construction. First, we explain the role of the mixture supermartingale $\bar{E}_{t}$ and why a single-parameter exponential martingale cannot yield a time-uniform bound without sacrificing optional stopping. Second, we provide a heuristic “continuous optimization” argument that motivates the geometric grid over $\lambda_{k}$ and the emergence of an iterated-logarithm term in the boundary.
@@
The mixture construction in Lemma 4.2 reconciles these goals by fixing a countable family of parameters a priori and combining the corresponding exponential supermartingales. This “method of mixtures” (also called pseudo-maximization) goes back to Robbins and Siegmund’s classical work on sequential boundary crossing and is developed systematically in the theory of self-normalized processes (see De la Pena et al. [2009]). Modern confidence-sequence constructions build on this principle in various forms, including explicit mixture and stitching frameworks such as the sub-$\psi$ approach in Howard et al. [2021], as well as betting- and e-process-based methods that aggregate families of nonnegative supermartingales; see Waudby-Smith and Ramdas [2024]. $\clubsuit$
@@
The mixture construction in Lemma 4.2 hinges on choosing a geometric grid of tuning parameters $\lambda_{k}=\mathrm{e}^{-k/2}$. Here we provide intuition for this choice by examining how the envelope $B_{k}(\Sigma_{t,\mathrm{eff}}^{2})$ is minimized, and how the geometry of the grid matches the natural scale of the problem. Let $v\coloneqq\Sigma_{t,\mathrm{eff}}^{2}$ and recall the upper envelope
$B_{k}(v):=\frac{\log(1/(\alpha w_{k}))}{\lambda_{k}}+2\lambda_{k}v,\qquad\lambda_{k}=\mathrm{e}^{-k/2},\quad w_{k}=\frac{6}{\pi^{2}(k+1)^{2}}.$
@@
$z^{\star}=\sqrt{\frac{2v}{D}}\ \implies\ \lambda^{\star}=\frac{1}{z^{\star}}=\sqrt{\frac{D}{2v}}.$
@@
The slowly varying term $\log(k+1)$ introduces an iterated-logarithm correction, where the $\log\log(\mathrm{e}+v)$ factor arises precisely from the slow growth of $\log(k+1)$ on this geometric grid. In particular, for a fixed time $t$, the ideal tuning satisfies $\lambda^{\star}(t)\asymp\Sigma_{t,\mathrm{eff}}^{-1}$. Since $\Sigma_{t,\mathrm{eff}}^{2}$ is unknown in advance and varies with the trajectory, a fixed geometric grid of the form $\lambda_{k}=\mathrm{e}^{-k/2}$ ensures that for every $t$ there exists a grid point $\lambda_{k}$ within a constant factor of the ideal choice $\lambda^{\star}(t)$.
@@
In short, the grid is chosen so that: (i) it spans the relevant scales of $\Sigma_{t,\mathrm{eff}}^{2}$ via exponentially spaced search points, and (ii) at least one $\lambda_{k}$ is guaranteed to be within a constant factor of the ideal minimizer, thereby achieving near-optimal behavior across all times $t$. This “scale-matching” intuition for geometric discretizations is common in mixture and stitching arguments for martingale concentration; see, for example, the heuristic discussion surrounding Theorem 1 in Howard et al. [2021] and the broader treatment of pseudo-maximization via mixtures in self-normalized processes De la Pena et al. [2009]. $\clubsuit$
4.3 Proof of Theorem 2.5
The proof of Theorem 2.5 follows the same overall structure as in the convex case, but the technical arguments differ in substantive ways due to the shift from optimality to stationarity and from convexity to smoothness assumptions. As the time-uniform concentration machinery is already developed in the convex setting, we adopt a more direct presentation here, emphasizing the distinct smoothness-based decomposition and the resulting martingale structure that drive the nonconvex analysis.
In contrast to the convex analysis, where the argument centers on the process $\bar{X}_{t}$, the nonconvex proof is organized around the cumulative martingale term
$\bar{Y}_{t}:=-\sum_{s=1}^{t}\eta_{s}\langle\nabla f(x_{s}),\xi_{s}\rangle,\qquad\text{with }\xi_{s}:=g_{s}-\nabla f(x_{s}),$ (32)
which captures the aggregated effect of stochastic gradient noise along the SGD trajectory. Using a smoothness-based descent inequality, we rewrite the weighted stationarity measure in terms of a descent contribution and the martingale term $\bar{Y}_{t}$. We then establish conditional sub-Gaussian control of $\bar{Y}_{t}$ and construct an associated exponential supermartingale. A time-uniform concentration bound for $\bar{Y}_{t}$, combined with a telescoping argument and normalization by $S_{t}$, yields an anytime-valid upper confidence sequence for the weighted stationarity measure $\bar{G}_{t}$.
We begin by deriving a decomposition of the weighted stationarity measure along the SGD trajectory. In contrast to the convex case, where progress is tracked through the distance process $Z_{t}$, the nonconvex analysis relies on the smoothness descent inequality. The following lemma expresses the cumulative weighted squared gradient norm as the sum of a smoothness-based descent term and a martingale error term induced by stochastic gradient noise.
Lemma 4.6 (Weighted stationarity decomposition).
Under Assumption 1.3, let $\{x_{s}\}$ be generated by the SGD recursion (1) with $\mathcal{X}=\mathbb{R}^{d}$. Then, for every $t\geq 1$,
$\sum_{s=1}^{t}\eta_{s}\|\nabla f(x_{s})\|^{2}\;\leq\;f(x_{1})+\frac{L}{2}\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}+\bar{Y}_{t},$ (33)
where $\bar{Y}_{t}$ is defined in (32).